eight times the sum of three consecutive odd integers is 210 more than 10 times the middle integer. Find the integers?
the answer is 13, 15, 17

Answer by mananth(16949)

eight times the sum of three consecutive odd integers is 210 more than 10 times the middle integer. Find the integers?

Let the numbers be x, x+2, x+4
8 times the sum of the consecutive odd numbers = 8(x+x+2+x+4)
10(x+2) +210
8(x+x+2+x+4) = 10(x+2) +210
24x+48=10x+20 +210
14x=230-48
14x=182
x=13
x+2=15
x+4=17
The numbers are 13,15,17